OpenStudy (anonymous):

Solve for x: (x - 9)^2

OpenStudy (anonymous):

x- 9 = 0

OpenStudy (anonymous):

18

OpenStudy (anonymous):

how did you get 18?

OpenStudy (anonymous):

x = 2x - 18 18 = 2x -x 18 = x

OpenStudy (anonymous):

so first you square the entire thing?

OpenStudy (anonymous):

oh that that's a square..

OpenStudy (anonymous):

yes sorry

OpenStudy (anonymous):

please write (x - 9)^2

OpenStudy (anonymous):

so is the first step to square it?

OpenStudy (anonymous):

then it would come out to x= x^2 - 81?

OpenStudy (anonymous):

hello?

OpenStudy (anonymous):

x = (x - 9)^2 x = x^2 - 18x + 81 0 = x^2 - 19x + 81 (now use quadratic formula to solve x) x = 6.46 or x = 12.54

OpenStudy (anonymous):

x- 9 = 0 => x = 9